# Object-level Geometric Structure Preserving for Natural Image Stitching

### v2

@@ 5 Experiments
Figure 3: APAP [58] and SPHP [32] see misalignment, and we delineated the indistinct portions using color-coded boxes. Autostitch [3], APAP [58], SPHP [32] and GSP [7] sees distortion. The convergence of the blue and red lines is essential, and we signify distortion by the intersection of these two lines. GES-GSP [14] successfully prevents distortion, but it undergoes misalignment. Our method addresses misalignment and distortion well.
@@
Evaluation metrics. We quantitatively assess the quality of our stitching results from two perspectives: distortion prevention and alignment. First, we employ the Mean Distorted Residuals (MDR) metric introduced in GES-GSP [14] to measure the degree of image distortion. In intuitive terms, if the points on the same side of the mesh were originally collinear and remain collinear after stitching, it implies that the stitching result has minimal distortion. Please refer to supplementary material of [14] for more details of MDR. Furthermore, we employ the Naturalness Image Quality Evaluator (NIQE) [38] metric to evaluate alignment performance. NIQE measures image clarity, and stitching results with misalignment will produce blurry areas, leading to worse NIQE scores. Compared to other image quality assessment methods like Structure Similarity Index Measure (SSIM) [49] and Peak Signal-to-Noise Ratio (PSNR), NIQE does not require a reference image. Since there is no strict ground truth for image stitching, we use NIQE to assess the quality of our stitching results. However, it is essential to emphasize that in the field of image stitching, there is no universally accepted objective evaluation method for assessing the quality of stitching results or the degree of distortion. Many practitioners still rely on subjective assessments based on human visual perception to gauge the quality of stitching outcomes.
Datasets. We utilize datasets commonly associated with previous image stitching algorithms, which include datasets sourced from AANAP [31], APAP [58], CAVE [25], DFW [30], DHW [15], GES-GSP [14], LPC [24], SEAGULL [33], REW [29], SVA[35] and SPHP [32]. We also collected our own dataset of 18 image pairs for image stitching. Thus the datasets used in experiments encompass a diverse range of settings, including indoor and outdoor environments, various regions around the world, urban and natural landscapes, among others.
@@
Baselines. GSP [7] is a great work in alignment. Based on GSP, GES-GSP [14] protects line and curve structures. Our method tries to protect global structure based on GSP. The ideas behind these three methods are closely related. UDIS++ [40] is most famous as a pioneering work in deep learning based image stitching. Since it doesn’t explicitly use feature points, we are unable to measure its quality with MDR here. We provide a detailed comparasion between OBJ-GSP and UDIS++ in the supplementary materials.
Implementation details. We evaluated the results of GSP [7] and GES-GSP [14] using their publicly available C++ code. Our stitching code is also implemented in C++. Our implementation of SAM includes two approaches. Drawing from the Segment Anything C++ Wrapper [12], we exported the SAM [28] encoder and decoder into Open Neural Network Exchange (ONNX) format and subsequently replicated SAM’s automatic mode within C++. To achieve the best possible stitching results, we also directly implemented the semantic segmentation component using the SAM’s publicly available Python code and utilized their semantic segmentation results. The stitching part of our experiments ran on the CPU, while the SAM modules were capable of running on both CPU and GPU.
@@
Quantitative results. Table 1 shows MDR and NIQE results on datasets used in other stitching algorithms and our own dataset. We outperform GSP and GES-GSP in both alignment and shape preservation. UDIS++ [40] is a good try in deep learning based image stitching, but the performance is still no better than traditional methods(GSP, GES-GSP and ours).
@@
Qualitative results. In the context of image stitching, human perception is primarily influenced by two key factors: alignment and distortion. Misalignment can result in undesirable artifacts such as ghosting and blurriness, while distortion can make the stitched image appear unnatural. Fig. 3 illustrates the stitching outcomes of six different methods, where we use straight lines and boxes to demonstrate the effects of alignment and distortion. Fig. 1 and 4 elucidates the reasons behind the superior performance of our OBJ-GSP method. With the assistance of semantic segmentation techniques, we place greater emphasis on preserving critical structures and ensuring holistic protection at the object-level for objects within the images. Please kindly refer to our supplementary material for more quanlitative results.
6 Ablation Studies
Lightweight SAM backbones. To assess the influence of semantic segmentation results on the stitching performance, we conducted a comparative analysis across three different backbones of the SAM [28] model, namely ViT-B, ViT-L, and ViT-H [13], with a progression from smaller to larger models. Larger models are inherently capable of capturing more fine-grained semantic details. The corresponding stitching results are depicted in Fig. 5. It is worth noting that the models based on ViT-B and ViT-L exhibit some blurriness and minor distortions. From the perspective of MDR, our model, under the three aforementioned backbone configurations, achieved improvements of 1.9%, 2.7%, and 3.6% over the baseline model GES-GSP [14], respectively.
@@ 6 Ablation Studies
Figure 5: GES-GSP and the proposed OBJ-GSP with three semantic segmentation backbones.
@@
Sampling strategies. Incorporating the SAM [28] with ViT-H [13] backbone, the utilization of triangular grid sampling yields superior results compared to the triangular sampling proposed in GES-GSP [14]. The triangular grid preserves the shape of lines but fails to maintain the overall geometric structure of the image. As illustrated in the Fig. 6, we have outlined the structural elements in the original image using red dashed lines and then superimposed them onto the combined results of triangular sampling and triangular grid sampling. Triangular grid sampling retains the positional relationships between the lines present in the original image.
Figure 6: From top to down: one of the input images, OBJ-GSP with triangle sampling and with triangle mesh sampling.
7 Discussion
Limitations: While OBJ-GSP has demonstrated state-of-the-art performance in image stitching by extracting object-level geometric structures with SAM and preserving them with triangle mesh sampling, it has several limitations:
1. Our method introduces a large semantic segmentation model into image stitching, resulting in higher computational costs. However, with the development of semantic segmentation techniques, lighter versions of SAM will emerge [53, 59], enhancing the speed of our work. According to our analysis, the effectiveness of geometric structure extraction significantly impacts the final results. Our method is constrained by the quality of SAM’s results. Smaller models like SAM [28] with Vit-B/L [13] do not perform as SAM with Vit-H. Fig. 5 demonstrates the differences. To stitch a pair of 800*600 images, SAM spends 25s on RTX2090 with 8-24G GPU memory, depending on the backbone. For C++ ONNX implementation, SAM VIT-B spends 1.5 min on CPU. Mesh optimization and image processing cost less than 4s on an Intel i5 CPU, almost the same as GES-GSP [14]. OBJ-GSP needs more computational resources and time than GES-GSP, but the stitching quality is also better. The time cost of SAM is larger than line detection methods in GES-GSP. The time used for triangle mesh optimization is almost the same as that in GES-GSP.
2. As shown in Fig. 7, the proposed method fails in cases of significant parallax. In the perspective of the left image, the corner of the building appears to be obtuse. However, based on the inference from the right image, the corner of the building should be a right angle. Consequently, the stitching algorithm becomes perplexed, unsure of how to preserve the shape of the house.
In only rare instances OBJ-GSP experiences failures due to the malfunction of SAM [28], such as in conditions characterized by inadequate illumination or an exceedingly sparse set of features. Please refer to the supplementary material for details.
Applications of OBJ-GSP: In many fields, there is a need for high-quality stitched images, even at the expense of long time costs and significant computational resources. In medical image processing [46, 48, 61], for instance, stitching multiple pathological slice images to reconstruct the entire tissue structure or organ’s three-dimensional model demands high precision and quality. These tasks typically necessitate precise alignment and seamless fusion, and can tolerate longer computational times to ensure accuracy. Similarly, in photography [36] or virtual tourism [44, 50] applications, stitching numerous high-resolution images is necessary to generate high-quality panoramic images. This holds true in the fields of remote sensing [56, 9, 60] and movie industry [37, 19] as well.
Post-processing in image stitching: Our method primarily addresses computing transformation matrices to achieve alignment and shape preservation. Post-processing techniques can be combined with our approach to achieve a more natural stitching effect. Blending [3, 54, 55] and seam-driven [16] methods can be used to further reduce blurring, while global straightening [15] can decrease distortion.
Broader impacts: It is important to acknowledge that we do not explicitly discuss broader impacts in the proposed OBJ-GSP image stitching algorithm, such as fairness or bias. Segment Anything Model (SAM) [28] has discussed its broader impacts regarding geographic and income representation, as well as Fairness in segmenting people. Further research into how our algorithm may interact with other aspects of image stitching is encouraged.
@@ 8 Conclusion
In this paper, we propose Object-level geometric structure preserving for natural image stitching (OBJ-GSP) algorithm, which stands as a novel approach to achieving natural and visually pleasing composite images. The existing state-of-the-art, GES-GSP [14], while effective, faced limitations in accurately capturing geometric structures and preserving spatial relationships between these structures. In response, OBJ-GSP leverages semantic segmentation techniques, specifically the Segment Anything Model (SAM) [28], to extract object-level geometric structures imbued with semantic information. The proposed OBJ-GSP method addresses the shortcomings of its predecessor by preserving structures that align more intuitively with human perception. It employs triangular mesh deformations within individual objects to safeguard not only individual geometric structures but also the overall shapes of objects within the images. Empirical evaluations across multiple image stitching datasets have unequivocally demonstrated that OBJ-GSP surpasses the performance of GES-GSP, establishing a new state-of-the-art benchmark in image stitching. In conclusion, the OBJ-GSP algorithm offers a promising direction for image stitching, enhancing the preservation of geometric structures and object-level details, ultimately leading to more natural and visually appealing composite images.
We would like to clearly show some experiment and dataset details in this supplementary material. In the supplementary material, we provide more qualitative results, and in depth discussion and results on why Segment Anything Model [28] helps the proposed OBJ-GSP image stitching model, and failures cases of SAM.
@@ 9 Ablation of EfficientSAM and mesh sampling
10 Dataset
We employed handheld smartphones to capture images. During the image acquisition process, we took care to minimize translational movement of the smartphone, primarily relying on rotation to adjust the framing. This approach was employed to ensure that the disparity between images remained manageable, preventing situations where the occlusion relationships between two images would be too dissimilar for successful image stitching. We amassed a total of 18 pairs of image sets, encompassing diverse scenes such as rooms, culinary creations, sculptures, gardens, rivers, ponds, industrial facilities, roads, and exteriors of buildings, among others. Additionally, we curated existing image stitching datasets [31, 58, 25, 30, 15, 14, 24, 33, 29, 35, 32] to supplement our data collection efforts. Ultimately, we constructed a dataset consisting of 124 groups of images, marking it as the largest dataset currently employed in image stitching endeavors. We intend to release this dataset to the public for further research and development.

### v4

@@ 4 Experiments
4.1 StitchBench
Previous work often collected a small number of images themselves and performed qualitative tests only. Meanwhile, they have different focuses, such as parallax between the foreground and background, sparse features in natural scenery, precise alignment and no distinct structures to preserve, and distinct line structures, without comprehensively evaluating models’ performance in a wide range of scenarios. To address the issue, we present the most extensive image stitching benchmark to date: StitchBench, which include 122 pairs of images from 12 works. We collect 18 pairs of images captured by cameras, in which the preservation of object structures is crucial. StitchBench also includes 7 sets of urban scenes captured by low-altitude drones, featuring tall buildings and requiring the assistance of segmentation models. To overcome our subjective preferences and the limited locations where we collected the images, we also collect test images used in previous state-of-the-art works, namely AANAP [23], APAP [45], CAVE [32], DFW [22], DHW [12], GES-GSP [11], LPC [17], SEAGULL [25], REW [21], SVA [26] and SPHP [24]. StitchBench is currently the most comprehensive stitching test dataset. An algorithm should demonstrate general applicability to perform well on all subsets of StitchBench: aligning well and preventing distortion naturally.
@@ 4.1 StitchBench
Evaluation metrics. We quantitatively assess the quality of our stitching results from two perspectives: distortion prevention and alignment. First, we employ the Mean Distorted Residuals (MDR) metric to measure the degree of image distortion. In intuitive terms, if the points on the same side of the mesh were originally collinear and remain collinear after stitching, it implies that the stitching result has minimal distortion. Furthermore, we employ the Naturalness Image Quality Evaluator (NIQE) [29] metric to evaluate alignment performance. We argue that NIQE is a more intitutive and better indicator of alignment than RMSE, SSIM and PSNR, as it measures image clarity, and stitching results with misalignment will produce blurry areas, leading to worse NIQE scores.
4.2 Baselines
@@
We compare with GSP [6] and GES-GSP [11]. UDIS [30] and UDIS++ [31] are famous works in applyig deep learning into image stitching. Since they do not explicitly use feature points, we are unable to measure its quality with MDR. We provide a detailed comparasion between OBJ-GSP and UDIS++ in the supplementary materials.
4.3 Results
@@
Quantitative results. Table 1 shows MDR and NIQE results on datasets used in other stitching algorithms and our own dataset. We outperform GSP and GES-GSP in both alignment and shape preservation. UDIS++ [31] is a good try in deep learning based image stitching, but the performance is still no better than ours.
Qualitative results.
@@
Fig. 1 and 2 elucidates the reasons behind the superior performance of our OBJ-GSP method. With the assistance of semantic segmentation techniques, we place greater emphasis on preserving critical structures and ensuring holistic protection at the object-level for objects within the images. Fig. 4 and 5 illustrates the stitching outcomes of six different methods, where we use straight lines and boxes to demonstrate the effects of alignment and distortion. Please kindly refer to our supplementary material for more quanlitative results.
@@
Figure 7: GES-GSP and the proposed OBJ-GSP with three Segment Anything Model backbones.
Figure 8: Triangle sampling preserves the shapes of lines on the ground, while triangle mesh sampling fails in preserving.
4.4 Low-Altitude Aerial Image Stitching
Image stitching requires meeting one of two conditions [2]: either the camera’s optical center remains stationary while the camera rotates, or the scene only consists of objects that are far from the camera. Existing stitching algorithms mainly address the issue of stitching when these conditions are slightly violated. For low-altitude aerial images, where the flight height of the aircraft is around 100 meters but the height of buildings is no less than 20-40 meters, the camera’s optical center moves significantly during drone shooting, thus completely failing to meet the two assumptions for stitching. Moreover, if the left and right walls of a building are captured in two separate shots, it would be a logical error to include both walls in a panorama (for a cube, at most three faces can be seen at a time, and it’s impossible to see two opposing faces simultaneously). For stitching low-altitude aerial images, we first use a semantic segmentation model to segment the roofs and walls. We then calculate the height of the roofs and orthographically project the buildings onto the ground plane before stitching. In this scenario, the semantic segmentation model is essential for the stitching process [3]. The stitching pipeline and result is in Fig. 6.
5 Ablation Studies and Discussions
5.1 Lightweight SAM backbones
To assess the influence of semantic segmentation results on the stitching performance, we conducted a comparative analysis across three different backbones of the SAM [20] model, namely ViT-B, ViT-L, and ViT-H [10], with a progression from smaller to larger models. Larger models are inherently capable of capturing more fine-grained semantic details. The corresponding stitching results are depicted in Fig. 7. It is worth noting that the models based on ViT-B and ViT-L exhibit some blurriness and minor distortions. From the perspective of MDR, our model, under the three aforementioned backbone configurations, achieved improvements of 1.9%, 2.7%, and 3.6% over the baseline model GES-GSP [11], respectively. The application of EfficientSAM [41], time consumption of OBJ-GSP, and its real-world applications are included in supplementary materials.
5.2 Sampling strategies
@@
Incorporating SAM [20] with ViT-H [10] backbone, triangular mesh sampling yields superior results compared to the triangular sampling proposed in GES-GSP [11]. The triangular mesh preserves the shape of lines but fails to maintain the overall geometric structure of the image. As illustrated in the Fig. 8, we outline the structural elements in the original image using red dashed lines and then superimposed them onto the combined results of triangular sampling and triangular grid sampling. Triangular grid sampling retains the positional relationships between the lines present in the original image.
@@ 6 Conclusion
In this paper, we propose OBJect-level Geometric Structure Preserving for natural image stitching (OBJ-GSP) algorithm, which stands as a novel approach to achieving natural and visually pleasing composite images. OBJ-GSP protects object shapes by first segmenting them out, and then preserve the structures with triangle meshes. We also demonstrate that semantic segmentation is necessary when it comes to low-altitude aerial image stitching. We collect new test image pairs in common scenes and aerial imaging, and choose images from previous works, to establish the most comprehensive image stitching benchmark by far: StitchBench. Detailed experiments with comprehensive baselines in StitchBench demonstrate the effectiveness of OBJ-GSP.
7 Extensive Discussions
7.1 Limitations
While OBJ-GSP has demonstrated state-of-the-art performance in image stitching by extracting object-level geometric structures with semantic segmentation and preserving them with triangle mesh sampling, it introduces a large semantic segmentation model into image stitching, resulting in higher computational costs. However, with the development of semantic segmentation techniques, lighter versions of SAM will emerge [41, 46], enhancing the speed of our work. According to our analysis, the effectiveness of geometric structure extraction significantly impacts the final results. Our method is constrained by the quality of SAM’s results. Smaller models like SAM [20] with Vit-B/L [10] do not perform as SAM with Vit-H. To stitch a pair of 800*600 images, SAM spends 25s on RTX2090 with 8-24G GPU memory, depending on the backbone. For C++ ONNX implementation, SAM VIT-B spends 1.5 min on CPU. Mesh optimization and image processing cost less than 4s on an Intel i5 CPU, almost the same as GES-GSP [11]. OBJ-GSP needs more computational resources and time than GES-GSP, but the stitching quality is also better. The time cost of SAM is larger than line detection methods in GES-GSP. The time used for triangle mesh optimization is almost the same as that in GES-GSP.
7.2 Applications of OBJ-GSP
In many fields, there is a need for high-quality stitched images, even at the expense of long time costs and significant computational resources. In medical image processing [36, 34, 48], for instance, stitching multiple pathological slice images to reconstruct the entire tissue structure or organ’s three-dimensional model demands high precision and quality. These tasks typically necessitate precise alignment and seamless fusion, and can tolerate longer computational times to ensure accuracy. Similarly, in photography [27] or virtual tourism [35, 38] applications, stitching numerous high-resolution images is necessary to generate high-quality panoramic images. This holds true in the fields of remote sensing [44, 7, 47] and movie industry [28, 14] as well. Currently, the emergence of faster and more accurate segmentation models, such as EfficientSAM [41], makes our method even more promising. We provide detailed comparisons in the supplementary materials.
7.3 Post-processing in image stitching
Our method primarily addresses computing transformation matrices to achieve alignment and shape preservation. Post-processing techniques can be combined with our approach to achieve a more natural stitching effect. Blending [2, 42, 43] and seam-driven [13] methods can be used to further reduce blurring, while global straightening [12] can decrease distortion.
7.4 Broader impacts
It is important to acknowledge that we do not explicitly discuss broader impacts in the proposed OBJ-GSP image stitching algorithm, such as fairness or bias. Segment Anything Model (SAM) [20] has discussed its broader impacts regarding geographic and income representation, as well as Fairness in segmenting people. Further research into how our algorithm may interact with other aspects of image stitching is encouraged.
@@ 8 Ablation of EfficientSAM and mesh sampling
9 StitchBench Metadata
We employed handheld smartphones to capture OBJ-GSP images. During the image acquisition process, we took care to minimize translational movement of the smartphone, primarily relying on rotation to adjust the framing. This approach was employed to ensure that the disparity between images remained manageable, preventing situations where the occlusion relationships between two images would be too dissimilar for successful image stitching. We amassed a total of 18 pairs of image sets, encompassing diverse scenes such as rooms, culinary creations, sculptures, gardens, rivers, ponds, industrial facilities, roads, and exteriors of buildings, among others.
We also collect 7 sets of aerial images, each consist of about 9 pairs of images of urban scenes. We fly drones at 100-120 meters, in urban areas where there are buildings, roads, trees, etc [3]. We collect images with a DJI Mavic Air 2 and the image size is set to be 3000 × 4000 Pixels. Cameras on the drone are kept vertical to the ground (bird view).
Additionally, we curate existing image stitching datasets [23, 45, 32, 22, 12, 11, 17, 25, 21, 26, 24] to supplement our data collection efforts. Ultimately, we constructed a dataset consisting of 122 groups of images, marking it as the largest dataset currently employed in image stitching endeavors. We release this dataset to the public for further research and development.
@@ 10 Reestablishing baselines
10.1 Implementation details
We evaluated the results of GSP [6] and GES-GSP [11] using their publicly available C++ code. Our stitching code is also implemented in C++. Our implementation of SAM includes two approaches. Drawing from the Segment Anything C++ Wrapper [9], we exported the SAM [20] encoder and decoder into Open Neural Network Exchange (ONNX) format and subsequently replicated SAM’s automatic mode within C++. To achieve the best possible stitching results, we also directly implemented the semantic segmentation component using the SAM’s publicly available Python code and utilized their semantic segmentation results. The stitching part of our experiments ran on the CPU, while the SAM modules were capable of running on both CPU and GPU.
10.2 Baselines
@@
11 Algorithm and Results of Low-Altitude Aerial Image Stitching
11.1 Why is Semantic Segmentation Necessary
For low-altitude drone aerial photography in urban areas, the drone’s flight altitude is low while the buildings are tall. Due to the significant distance difference between the drone camera and the buildings, the transformation matrices for buildings and rooftops differ from those for the ground in different images. Direct stitching can lead to a lot of ghosting and unnatural, distorted building structures. Additionally, selective information must be discarded during low-altitude aerial stitching in urban areas. If buildings are simply considered rectangular prisms, and the left and right walls are captured in two separate shots, it is impossible to retain both sides in the stitched image (as a person cannot see both opposing sides of a rectangular prism simultaneously). Therefore, this paper proposes first using a segmentation model to identify buildings and walls in the scene. The walls are removed from the images, and the ratio of the building height to the drone height is calculated based on the different transformation relationships for the ground and buildings. The buildings are then projected onto the ground plane before stitching. The importance of segmentation in aerial image stitching is shown in Fig. 10
Figure 10: In these images, drones fly at about 100m height. Top: segmentation and orthographic projection is not necessary when everything in the images are relative low. In top left image, the captures scene is almost a plain, and the towers is about 6 meters tall. The buildings are two-floor villas in the bottom left image, and no more than 8 meters tall. Middle: The building in the middle left images are 6 floors, and about 20 meters tall. In middle right images, the buildings are about 15 meters tall. The tall buildings can not be stitched by traditional stitching methods. Bottom: OBJ-GSP can stitch the images by segmenting roofs and walls out, transform roofs with orthographic projection,and then stitch.
11.2 OBJ-GSP in Aerial Image Stitching
In aerial images of low altitude, there are two planes of interest: ground $P^{g}$ and roofs $P^{r}_{i}$, where $i=1,2,...N$ and $N$ is the number of roofs. Semantic segmentation models are adopted to segment roofs and walls, with the remaining pixels regarded as ground. We mask out walls, and then orthographically project roofs to grounds. For correctly matched feature points ${f}_{r1}$ and ${f}_{r2}$on $P^{r}$, and ${f}_{g1}$ and ${f}_{g2}$ on $P^{g}$, we aim to find transformation matrix $H_{o}$ to project roof to ground before stitching. After projection, the images can be stitched with a global transformation matrix, which is also the transformation between grounds, $H_{g}$.
$\begin{cases}f_{g2}=H_{g}f_{g1}\\ H_{oi}f_{r2}={H_{g}}H_{oi}f_{r1}\end{cases}\text{for}\ i\ \text{in}1\ \text{to}\ N\text{,}$ (11)
where $f$ is in homogeneous coordinates $\left[\begin{matrix}x&y&1\\ \end{matrix}\right]^{T}$. Let the height of roof $P^{r}_{i}$ to the ground be $h_{ri}$, and height of drone be $h_{d}$, for each pixel $[x,y,1]$ on the roof, the orthographic projection transformation is
$\left[\begin{matrix}x_{g}\\ y_{g}\\ 1\\ \end{matrix}\right]=\ \left[\begin{matrix}1-\frac{h_{ri}}{h_{d}}&0&\frac{h_{ri}}{h_{d}}a\\ 0&1-\frac{h_{ri}}{h_{d}}&\frac{h_{ri}}{h_{d}}b\\ 0&0&1\\ \end{matrix}\right]\ \left[\begin{matrix}x_{r}\\ y_{r}\\ 1\\ \end{matrix}\right]\text{for}\ i\ \text{in}1\ \text{to}\ N\text{,}$ (12)
where $a$ and $b$ are half of pixel width and height of the image. If the transformation matrix is in the form of a homograph matrix, and the number of feature point matches is $M_{i}$ on $P^{r}_{i}$, Eq. 11 expands to $2M_{i}$ mutually independent quadratic equations with the unknown $\frac{h_{ri}}{h_{d}}$. This over determined equation can be solved by methods such as Newton iteration. After solving $\frac{h_{ri}}{h_{d}}$, the orthographic projection map can be generated by simply transforming $P^{r}_{i}$ to the ground one by one. The transformation matrix is the affine transformation matrix, which can be solved similarly for the orthographic projection matrix.
11.3 Segmentation Models and Aerial Segmentation Datasets Used
We finetune Grounded SAM [33] on low-altitude drone datasets where roof and wall are annotated (Varied Drone Dataset [4] and ICG Drone Dataset [16]).